How do I solve abs(x-3)=x-1 graphically?
Answer by jim_thompson5910(35256)

Graph both y = |x-3| (in green) and y = x-1 (in blue) to get this



You can see that the two graphs intersect at the point (2,1). So the solution to |x-3| = x-1 is x = 2